OpenStudy (anonymous):

2x^6+8x^2=9 rational roots

OpenStudy (anonymous):

\[\pm1 \pm3 \pm9\]

OpenStudy (anonymous):

\[\pm1 \pm2 \pm4 \pm 8\]

OpenStudy (anonymous):

\[\pm1 \pm3 \pm9 \pm1/2 \pm3/2\pm9/2\]

OpenStudy (anonymous):

\[\pm1 \pm2 \pm4 \pm8 \pm1/2 \pm1/4 \pm1/8\]

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

got to find the constant term

OpenStudy (anonymous):

and leading coefficient

OpenStudy (anonymous):

form p/q

OpenStudy (jdoe0001):

ohhh hmmm

OpenStudy (jdoe0001):

right.... so.... hmm ok... so that should be simple just all the possible combinations?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

rational root theorem

OpenStudy (anonymous):

http://www.mathwords.com/r/rational_root_theorem.htm

OpenStudy (jdoe0001):

\(\large { 2x^6+8x^2=9 \implies \begin{array}{llll} 2x^6+8x^2-&9=0\\ \uparrow &\uparrow \\ p&q \end{array} }\)

OpenStudy (jdoe0001):

so just get the factors for each set the factors for "p" over the factors "q" keep in mind that "1" is also a factor

OpenStudy (anonymous):

1 3 9

OpenStudy (jdoe0001):

actaully

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

\(\large { 2x^6+8x^2=9 \implies \begin{array}{llll} 2x^6+8x^2-&9=0\\ \uparrow &\uparrow \\ q&p \end{array} }\)

OpenStudy (anonymous):

only way I see it as 1 3 9

OpenStudy (jdoe0001):

so "p factors are 3, 3, 1 and 9 and "q" factors are 2, 1 only so p/q factors will be \(\bf \pm\cfrac{3,3,1,9}{2,1}\)

OpenStudy (jdoe0001):

\(\large { \pm\cfrac{3,3,1,9}{2,1}\implies \begin{cases} \pm \frac{3}{2}\\ \pm \frac{3}{1}\\ \pm \frac{1}{2}\\ \pm \frac{1}{1}\\ \pm \frac{9}{2}\\ \pm \frac{9}{1}\\ \end{cases} }\)

OpenStudy (anonymous):

so it is C

OpenStudy (jdoe0001):

so.. .that'd be p/q

OpenStudy (anonymous):

c seems like a reasonable answer

OpenStudy (jdoe0001):

dunno what C is, but anyhow... that's the p/q if C shows that, then that's the one keep in mind that the choices are just guides they should match your answer, not the other way around